OpenStudy (anonymous):

\[\frac{ \sqrt{5\sqrt{5}} }{ \sqrt[8]{25} }\]

OpenStudy (anonymous):

\[\frac{ \sqrt{5\sqrt{5}} }{ \sqrt[8]{25} }\]

OpenStudy (unklerhaukus):

What is 25 in prime factors?

OpenStudy (anonymous):

\[\sqrt{5\sqrt{5}}\]

OpenStudy (anonymous):

= \[\sqrt{5} \times \sqrt{5} ?\]

OpenStudy (unklerhaukus):

nope, you can't do that, the large radical is over the √5 too

OpenStudy (unklerhaukus):

but you can have \[\sqrt{5\sqrt 5}=\sqrt5\sqrt{\sqrt5}\]

OpenStudy (unklerhaukus):

If you convert those radicals into indices \[\large \sqrt[q]x^p= x^{p/q}\] \[\sqrt{\sqrt5}=({5^{1/2}})^{1/2}=5^{\tfrac12\times\tfrac12}\]

OpenStudy (anonymous):

So \[\sqrt{5\sqrt{5}} = 5_{4}^{1} ?\]

OpenStudy (unklerhaukus):

yes, so what is \[\sqrt5\sqrt{5\sqrt5}\]

OpenStudy (unklerhaukus):

remember that \[\large a^na^m=a^{n+m}\]

OpenStudy (anonymous):

\[\sqrt[?]{25} = 5_{8}^{2}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

* \[\sqrt[8]{25}\]

OpenStudy (anonymous):

So \[5_{4}^{1} / 5_{8}^{2} = ?\]

OpenStudy (unklerhaukus):

wait a minute,

OpenStudy (unklerhaukus):

\[\sqrt{5\sqrt 5}=\sqrt5\sqrt{\sqrt5}\] \[\qquad\qquad\qquad\sqrt{\sqrt{5}} = 5^{1/4}\] \[\sqrt{5\sqrt 5}=\sqrt5\times 5^{1/4}\\ \qquad\quad=\]

OpenStudy (unklerhaukus):

\[\sqrt[8]{25} = 5^{2/8}\] is right by the way.

OpenStudy (anonymous):

\[\sqrt{5} \times 5_{4}^{1} = ?\]

OpenStudy (unklerhaukus):

yeah so use \[\sqrt 5=\sqrt[2] 5 =5^{1/2}\]

OpenStudy (anonymous):

so \[\sqrt{5\sqrt{5}} = 5_{4}^{1} rigth?\]

OpenStudy (anonymous):

2/8 *

OpenStudy (unklerhaukus):

nope \[\sqrt{5\sqrt{5}} = (5\times(5)^{1/2})^{1/2}=\]

OpenStudy (anonymous):

( 5 x \[5_{4}^{1}\] ) =

OpenStudy (unklerhaukus):